How many gallons of a 20% antifreeze solution and a 15% antifreeze solution must be mixed to obtain 30 gal of a 19% antifreeze solution? Round to the nearest gallon, if necessary.
Answer by Boreal(15235)

x gal of 20%
30-x gal of 15%
30 gal of 19%=5.7 gal pure antifreeze
the sum of the other two are .20x+0.15(30-x)=.20x+4.5-0.15x
so 0.05x+4.5=5.7
0.05x=1.2
x=24 gal of 20%. (4.8 gal pure)
30-x=6 gal of 15% (0.9 gal pure)
Check 19 is 1/5 of the way from 20 to 15, and 1/5 of 30 is 6 gal of the 15%. It is weighted much more towards the 20% antifreeze.